### v2

# Generalized Animal Imitator: Agile Locomotion with Versatile Motion Prior

Abstract
The agility of animals, particularly in complex activities such as running, turning, jumping, and backflipping, stands as an exemplar for robotic system design. Transferring this suite of behaviors to legged robotic systems introduces essential inquiries: How can a robot be trained to learn multiple locomotion behaviors simultaneously? How can the robot execute these tasks with a smooth transition? How to integrate these skills for wide-range applications? This paper introduces the Versatile Instructable Motion prior (VIM) – a Reinforcement Learning framework designed to incorporate a range of agile locomotion tasks suitable for advanced robotic applications. Our framework enables legged robots to learn diverse agile low-level skills by imitating animal motions and manually designed motions. Our Functionality reward guides the robot’s ability to adopt varied skills, and our Stylization reward ensures that robot motions align with reference motions. Our evaluations of the VIM framework span both simulation environments and real-world deployment. To the best of our knowledge, this is the first work that allows a robot to concurrently learn diverse agile locomotion skills using a single learning-based controller in the real world. Further details and supportive media can be found at our project site: https://rchalyang.github.io/VIM
Figure 1: Real-Robot Trajectory. Our robot demonstrates diverse agile locomotion skills, including running, jumping, and back-flipping in diverse real-world scenarios using a single motion prior and without any fine-tuning.
*Equal contribution
I Introduction
Researchers have been studying for years equipping legged robots with agility comparable to that of natural quadrupeds. Picture a golden retriever gracefully maneuvering in a park: darting, leaping over obstacles, and pursuing a thrown ball. These tasks, effortlessly performed by many animals, remain challenging for contemporary legged robots. To accomplish such tasks, robots need not only master individual agile locomotion skills like running and jumping but also the capacity to adaptively select and configure these skills based on sensory inputs. We regard this kind of complicated task requiring highly agile locomotion skills as advanced parkour for legged robots. The inherent ability of quadrupeds to smoothly execute diverse locomotion skills across varied tasks inspires our pursuit of a control system with a general locomotion motion prior that includes these skills. In this direction, we introduce a novel RL framework, Versatile Instructable Motion prior (VIM) aiming to endow legged robots with a spectrum of reusable agile locomotion skills by integrating existing agile locomotion knowledge.
Historically, agile gaits[4, 36, 37] for legged robots have been sculpted using model-based or optimization methods. While promising, these methods demand significant engineering input and precise state estimation. Learning-based controllers enable robots to walk or run while addressing these limitations, although they still fall behind in agility. Imitation-based controllers are also proposed to learn from motion sequences from animals [40] or optimization methods [16]. Research on incorporating visual observations [1, 57, 56, 24, 60, 27, 62] or elevation maps [33, 44] further enables legged robots to traverse complex terrain like stones. Despite encouraging results, most of these works focus on building a single controller from scratch, even though much of the learned locomotion skills could be shared across tasks. Recent works start building a reconfigurable low-level motion prior [46, 23, 5, 42, 26, 22] for downstream applications. However, previous methods failed to make the best use of existing skills to learn diverse locomotion skills with high agility.
Figure 2: Learning Agile Locomotion Skills from Reference Motions: Our system learns a single instructable motion prior, from a diverse reference motion dataset and makes the learned skills available for solving downstream tasks.
In this work, we focus on building low-level motion prior to utilize existing locomotion skills in nature and previous optimization methods, and learn multiple highly agile locomotion skills simultaneously, as in Figure 2. Even though we cannot fully comprehend the agility of animals and lack a unified framework for model-based controls, we recognize that motion sequences offer a consistent representation of diverse agile locomotion skills. Our motion prior extracts and assimilates a range of locomotion skills from reference motions, effectively mirroring their dynamics. These references comprise motion capture (mocap) sequences from quadrupeds, synthetic motion sequences complementing mocap data, and optimized motion trajectories. Throughout the training phase, we translate varied reference motion clips into a unified latent command space, guiding the motion prior to recreate locomotion dynamics based on these latent commands and the robot’s inherent state.
For legged robots, we define a locomotion skill as the ability of the robot to produce a specific trajectory. To break down the intricacies of movement, we classify it into two primary aspects: Functionality and Style. Functionality pertains to the fundamental movement objectives a robot aims to achieve, such as advancing forward at a predefined speed. Style, in contrast, focuses on the specific mechanics of how a robot accomplishes a task. To illustrate, two robots might be programmed to progress at an identical speed, but the intricacies of their movement—like step size or frequency—might differ considerably. Simultaneously instructing a robot in both these domains is nontrivial [55]. Drawing inspiration from how humans learn complicated tasks, especially in fields demanding physical prowess like athletics, we identify three core feedback modalities: objective performance metrics, qualitative assessments, and detailed kinematic guidance. Adopting this structured feedback approach, our robot starts with mastering the basic functional objective and subsequently turns into refining the detailed locomotion gaits.
By incorporating diverse reference motions and our reward design, our Versatile Instructable Motion prior (VIM) learns diverse agile locomotion skills and makes them available for intricate downstream tasks. With our VIM, we enable legged robots to perform advanced robotics parkour in the real world. We also evaluate our method in the simulation and real world, as Figure 1. Our method significantly outperforms baselines in terms of final performance and sample efficiency.
II Related Work
Blind Legged Locomotion: Classical legged locomotion controllers [19, 59, 52, 34, 4, 43] based on model-based methods [8, 17, 10, 12, 3, 20, 49] and trajectory optimization [7, 37] have shown promising results in diverse tasks with high levels of agility. Nonetheless, these methods normally come with considerable engineering work for the specific task, high computation requirements during deployment, or fragility to complex dynamics. Learning-based methods [1, 29, 33, 47, 9, 15, 31] controllers are proposed to offer robust and lightweight controllers for deployment at the cost of offline computation. Peng et al [14] developed a controller producing non-agile life-like gaits by imitating animals. Though previous works offer robust or agile locomotion controllers across complex environments, these works focus on finishing a single task at a time without reusing previous experience. Peng et al [41] leverage reference motions as prior knowledge when directly addressing specific tasks. Li et al [30] obtain an agile locomotion skill from a single partial rough demonstration including only robot root trajectory. Smith et al. [48] utilize existing locomotion skills to solve specific downstream tasks. Vollenweider et al. [54] utilize multiple AMP [41] to develop a controller to solve a fixed task set. In this paper, our motion prior captures diverse agile locomotion skills from reference motions generated by trajectory optimization and provides them for intricate future downstream tasks.
Motion Priors: Due to the low sample efficiency and considerable effort required for reward engineering of RL, low-level skill pre-training has drawn growing attention in recent years. Singh et al [46] utilize a flow-based model to build an actionable motion prior with motion sequences generated by scripts. More recent works [23, 5, 42, 26, 13, 22] focus on building low-level motion prior for downstream tasks but fail to include diverse highly agile locomotion skills. Peng et al [42] develop a simulation-based low-level motion prior entirely through unsupervised methods, yet they do not assure the acquisition of agile skills from the reference motion dataset. In this work, we build motion prior with reference motions consisting of mocap sequences, synthesized motion sequences, and trajectories from optimization methods and learn multiple highly agile locomotion skills with a single controller.
III Learn Versatile Instructable Motion Prior
We present the Versatile Instructable Motion prior (VIM), depicted in Figure 3, designed to acquire a wide range of agile locomotion skills concurrently from multiple reference motions. The development of our motion prior involves three stages: assembling a comprehensive dataset of reference motions sourced from diverse origins, crafting a motion prior that processes varying reference motions and the robot’s proprioceptive feedback to generate motor commands, and finally, utilizing an imitation-based reward mechanism to effectively train this motion prior.
III-A Reference motion dataset
TABLE I: Reference Motion Dataset: Our reference motions (11 reference motions in total) come from motion capture of animal motion, trajectory optimization method, and synthesized data with generative model whose length ranges from 32 to 500. During training, we repeat the reference motions cyclically to fit the length of the episode.
Skill Name Walk (Mocap) Trot (Mocap) Jump while Running (Mocap) Right Turn (Mocap) Left Turn (Mocap) Pace (Mocap)
Motion length 500 32 500 38 45 38
Skill Name Jump Forward (Synthetic) Left Turn (Synthetic) Backflip (Optimization) Jump Forward (Optimization) Canter (Mocap)
Motion length 500 500 129 120 64
Our primary objective is to curate a skill set for the robot that covers diverse functions and agility levels, equipping it to handle complex downstream tasks. Our dataset encompasses reference motions for locomotion skills, including but not limited to canter, pace, walk, trot, turns, backflips, and various jumps. These reference motions are derived from: (a) Mocap data of quadrupeds, specifically a subset from previous work [61], despite its inherent challenges like noise due to the unpredictability of animal behavior; (b) synthesized motions generated using a generative model [61], aimed at enhancing dataset diversity by capturing challenging locomotion actions; (c) motions crafted through trajectory optimization methods.
To address the methodology disparities between the real animals and our robot, we retarget both mocap and synthesized sequences to our robot as Peng et al. [40]. While mocap and synthesized motions offer extensive data, not all sequences may be practically achievable by the robot. Thus, our dataset is supplemented with motion sequences from trajectory optimization, emphasizing intricate moves like jumps and backflips. The comprehensive reference motion list can be found in Table I. Each trajectory in our dataset, represented as $(s^{\text{ref}}_{0},s^{\text{ref}}_{1},\cdots,s^{\text{ref}}_{T})$, where $s^{\text{ref}}_{i}$ is the robot state at $i$th timestep in the reference motion, and $T$ is the length of the reference motion. For every frame of the reference motion, it includes the following information: 1) robot pose in the world frame 2) robot velocity in the robot frame 3) foot position in the robot frame 4) foot height over the ground 5) robot joint angle, and 6) robot joint velocity. Note that reference motions do not include specific motor commands which are absent in the captured and synthesized data. For information that is hard to obtain in the real world, like robot pose and robot velocity, we only use them in simulation for reward calculation during training, our policy does not require that information in the real world. We denote the dataset as $\mathcal{D}=\{(s^{\text{ref}}_{0},s^{\text{ref}}_{1},\cdots,s^{\text{ref}}_{T})_{i}\}_{i=1}^{N}$, where $N$ is the total number of reference motions we have in the dataset.
III-B Motion Prior Structure
Figure 3: Versatile Instructable Motion prior (VIM): Our instructable motion prior consists of a reference motion encoder and a low-level policy. Our reference motion encoder maps reference motions into latent skill space indicating target robot pose and low-level policy output motor command. $V_{low}$ is the low-level critic for RL training. Reward Design: Our reward includes Functionality reward and Style reward. Our Functionality reward encourages the robot to track the root pose trajectory of the reference motion, and Style reward encourages the robot to track the joint motion of the reference motion.
Our motion prior consists of a reference motion encoder and a low-level policy. Reference motion encoder maps varying reference motions into a condensed latent skill space, and low-level policy utilizes our imitation reward and reproduces the robot motion given a latent command.
Reference motion encoder: Our reference motion encoder $\mathbb{E}_{\text{ref}}(\cdot)$ maps segments of reference motion to latent commands in a latent skill space that outlines the robot’s prospective movement. These segments are expressed as $\hat{s}^{\text{ref}}_{t}=\{s^{\text{ref}}_{t+1},s^{\text{ref}}_{t+2},s^{\text{ref}}_{t+5},s^{\text{ref}}_{t+10},s^{\text{ref}}_{t+30}\}$. Specifically, we choose $s^{\text{ref}}_{t+1},s^{\text{ref}}_{t+2},s^{\text{ref}}_{t+5}$ to provide immediate desired future motion, $s^{\text{ref}}_{t+10},s^{\text{ref}}_{t+30}$ to provide desired motion over a longer time-span. We model the latent command as a Gaussian distribution $\mathcal{N}(\mathbb{E}_{\text{ref}}^{\mu}(\hat{s}^{\text{ref}}_{t}),\mathbb{E}_{\text{ref}}^{\sigma}(\hat{s}^{\text{ref}}_{t}))$ from which we draw a sample at each interval to guide the low-level policy.
To maintain a temporal-consistent latent skill space, our training integrates an information bottleneck [51, 2] objective $L_{\text{AR}}$, where the prior follows an auto-regressive model [6]. Specifically, given the sampled latent command for the previous time step $z_{t-1}$, we minimize the KL divergence between the current latent Gaussian distribution and a Gaussian prior parameterized by $z_{t-1}$,
$L_{\text{AR}}(\hat{s}^{\text{ref}}_{t},z_{t-1})=\beta\text{KL}\left(\mathcal{N}(\mu_{t},\sigma_{t}^{2})\parallel\mathcal{N}(\alpha z_{t-1},(1-\alpha^{2})I)\right),$
where $\alpha=0.95$ is the scalar controlling the effect of correlation, $\beta$ is the coefficient balancing regularization.
Low-level policy training: Our low-level policy $\pi_{\text{low}}$ takes latent command $z_{t}$ representing the desired robot pose and robot’s current proprioceptive state $s_{t}$ as input, and outputs actual motor commands $a_{t}$ for the robot, where $s_{t}$ is encoded with a proprioception encoder $\mathbb{E}_{\text{prop}}$. We train low-level policy and reference motion encoder using PPO [45] in an end-to-end manner. Additionally, we introduce a motion embedding for the critic ($V_{low}$ in Figure 3) to distinguish diverse reference motions. Episodes initiate with randomized starting time steps from the dataset to avert overfitting and conclude when the root pose tracking error escalates beyond an acceptable range or the episode length is beyond the maximum length.
III-C Imitation Reward for Functionality and Style
Given the formulation of our motion prior, the robot learns diverse agile locomotion skills with our imitation reward and reward scheduling mechanics. Our reward offers consistent guidance, ensuring the robot captures both the functionality and style inherent to the reference motion.
Learning Skill Functionality: To mirror the functionality of the reference motion, we translate the root pose discrepancy between agent trajectories and reference motion into a reward. The functionality reward $r_{\text{func}}$ is subdivided into tracking rewards for robot root position $r_{\text{func}}^{\text{pos}}$ and orientation $r_{\text{func}}^{\text{ori}}$. Recognizing the distinct importance of vertical movement in agile tasks, the root position tracking is further split into rewards for vertical $r_{\text{func}}^{\text{pos-z}}$ and horizontal movements $r_{\text{func}}^{\text{pos-xy}}$.
$r_{\text{func}}(s_{t},\hat{s}_{t}^{\text{ref}})=w_{{\text{func}}}^{\text{ori}}*r_{\text{func}}^{\text{ori}}+w_{\text{func}}^{\text{pos-xy}}*r_{\text{func}}^{\text{pos-xy}}+w_{\text{func}}^{\text{pos-z}}*r_{\text{func}}^{\text{pos-z}}$
The specific formulation of our functionality rewards is provided as follows, which is similar to previous work[40].
$r_{\text{func}}^{\text{ori}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-10\left\lVert\hat{\mathbf{q}}_{t}^{\text{root}}-\mathbf{q}_{t}^{\text{root}}\right\rVert^{2}\right)$
$r_{\text{func}}^{\text{pos-xy}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-20\left\lVert\hat{\mathbf{x}}_{t}^{\text{root-xy}}-\mathbf{x}_{t}^{\text{root-xy}}\right\rVert^{2}\right)$
$r_{\text{func}}^{\text{pos-z}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-80\left\lVert\hat{\mathbf{x}}_{t}^{\text{root-z}}-\mathbf{x}_{t}^{\text{root-z}}\right\rVert^{2}\right)$
where $\mathbf{q},\hat{\mathbf{q}}$ and $\mathbf{x},\hat{\mathbf{x}}$ denote the root orientation and position from both the robot and reference motion, respectively. Notably, in contrast to previous work [40], we allocate a greater emphasis on root height in our reward, crucial for mastering agile locomotion skills such as backflips and jumps.
Learning Skill Style: Capturing the style of a reference motion, in addition to its functionality, expands the application of the locomotion skills by meeting criteria such as energy efficiency, robot safety, and facilitating human-robot interaction. Drawing inspiration from how humans learn - starting by emulating the broader style before focusing on intricate joint movements - our robot first mimics the broader locomotion style with an adversarial style reward and later refines its technique with a joint angle tracking reward.
Adversarial Stylization Reward: To quickly capture the style of the locomotion skill, we train distinct discriminators $D_{i},\ i=1..N$ for all $N$ reference motions separately to distinguish robot transitions from the transition of the specific reference motion [41, 54] and use the output to provide high-level feedback to the agent. Specifically, our discriminator is trained with the following objective:
$\mathop{\mathrm{argmin}}_{D_{i}}$ $\mathop{\mathbb{E}}_{d_{i}^{\mathcal{M}}(s_{t},s_{t+1})}\left(D_{i}(s_{t},s_{t+1})-1\right)^{2}$
$+$ $\mathop{\mathbb{E}}_{d_{i}^{\pi}(s_{t},s_{t+1})}\left(D_{i}(s_{t},s_{t+1})+1\right)^{2}$
where $d_{i}^{\mathcal{M}}(s_{t},s_{t+1})$ and $d_{i}^{\pi}(s_{t},s_{t+1})$ denote the state transition pair distribution sampled from the dataset, and the state transition pair distribution generated by the policy for $i$th reference motion respectively.
For each reference motion, the likelihood from the discriminator is then converted to a reward with:
$r_{\text{style}}^{\text{adv}}(s_{t},s_{t+1})=1-\frac{1}{4}*\left(1-D(s_{t},s_{t+1})\right)^{2}$
Initially, our adversarial stylization reward provides dense reward and enables the robot to learn a credible gait, but it can not provide more detailed instructions as the training proceeds, which leads to mode collapse and unstable training.
Joint Angle Tracking Reward: On the contrary, joint angle tracking reward [39, 5] provides sparse but stable instruction for the robot to mimic the gait of reference motion. Similar to our root pose tracking reward, our joint angle tracking reward has the following formulation:
$r_{\text{style}}^{\text{joint}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-5\sum_{j\in\mathrm{joints}}\left\lVert\hat{\mathbf{q}}_{t}^{j}-\mathbf{q}_{t}^{j}\right\rVert^{2}\right)$
$+\exp\left(-20\sum_{f\in\mathrm{feet}}\left\lVert\hat{\mathbf{e}}_{t}^{f}-\mathbf{e}_{t}^{f}\right\rVert^{2}\right)$
$+\exp\left(-20\sum_{f\in\mathrm{feet}}\left\lVert\hat{\mathbf{h}}_{t}^{f}-\mathbf{h}_{t}^{f}\right\rVert^{2}\right)$
where $\mathbf{q}_{t}^{j},\hat{\mathbf{q}}_{t}^{j}$ are the joint angle of robot and reference motion, $\mathbf{e}_{t}^{f},\hat{\mathbf{e}}_{t}^{f}$ are the end-effector positions of robot and reference motion, $\mathbf{h}_{t}^{f},\hat{\mathbf{h}}_{t}^{f}$ are the end-effector height of robot and reference motion.
Stylization Reward Scheduling: To learn the style quickly and stably, we propose to use both adversarial stylization reward and joint angle tracking reward with a balanced scheduling mechanism. Considering the discriminator as a "coach", we utilize the mean adversarial reward as an indication of how the coach is satisfied with the current performance. When it’s not satisfied with the current performance of the robot, it provides detailed instructions for the robot to learn. Specifically, our stylization reward follows:
$r_{\text{style}}(s_{t},\hat{s}_{t}^{\text{ref}})$ $=w_{{\text{style}}}^{\text{adv}}*r_{\text{style}}^{\text{adv}}+w_{\text{style}}^{\text{joint}}*r_{\text{style}}^{\text{joint}}$
$+w_{{\text{style}}}^{\text{adv}}*(1-\displaystyle\mathop{\mathbb{E}}_{s_{t}\in S}(r_{\text{style}}^{\text{adv}}(s_{t},s_{t+1})))*r_{\text{style}}^{\text{joint}}$
With the given formulation, our stylization reward provided dense rewards during the beginning of training, enabling the robot to quickly catch the essence of different agile locomotion skills. Our stylization reward also provides detailed instruction as the training proceeds, enabling the robot to refine its gait and lead to more stable training.
III-D Solving Downstream Tasks with Motion Prior:
Figure 4: Solving High-level Tasks with Our Motion Prior. Our high-level policy takes the high-level task information and the proprioception as input and output high-level latent command for low-level policy.
For intricate tasks like jumping over gaps, addressing them from scratch presents challenges including acquiring necessary agile locomotion skills, such as jumping and rapid running within limited interactions, and the intensive engineering needed to harmonize the reward for top-tier tasks while regularizing the robot’s motion. With a low-level motion prior, robots can instantly harness existing skills encapsulated within the prior and channel their efforts into high-level strategizing. For each distinct downstream task, we train a high-level policy $\pi_{\text{high}}$ (As shown in Figure 4) takes the high-level observation $\mathbf{o}_{\text{high}}$, and proprioception of the robot and outputs latent command for low-level motion prior to utilize the existing agile locomotion skills: $a_{t}=\pi_{\text{low}}(\pi_{\text{high}(\mathbf{o}_{\text{high}},s_{t}),\mathbf{E}_{\text{prop}}(s_{t})})$.
III-E Implementation Details:
Observation Space: Our low-level observation includes joint angles, joint velocities, gravity vector in the robot frame, and the previously executed actions.
Action Space: Our action space is defined as target joint angles for 12 joints of our robot. The target joint angles are converted to torque command with PD controller where KP=40, KD=1.0. Our control frequency in simulation and real-world is 25Hz.
Reference Encoder $E_{ref}$: Our reference encoder is a two-layer MLP with $[256]$ hidden units, mapping the reference motion segment into a $64$ dimensional latent distribution.
Proprioception Encoder $E_{prop}$: Our reference encoder is a two-layer MLP with $[256]$ hidden units, mapping proprioception information into a $64$-dim robot state feature.
Low-level Policy $\pi_{low}$ and Value Network $V_{low}$: Our low-level policy is a three-layer MLP with $[256,128]$ hidden units, mapping the robot state feature and latent command to $12$-dim robot target joint angles. Our low-level value network shares the same structure with an additional $64$-dim motion embedding as input, and output 1-dim value for RL training.
High-level Policy $\pi_{high}$ and Value Network $V_{high}$: Our high-level is formulated as a three-layer MLP with $[256,128]$ hidden units, mapping proprioception information and high-level task information to high-level latent command for low-level motion prior. Our high-level value network shares the same structure. High-level task information depends on specific task.
Low-level training termination: As stated in Sec 3, we terminate the episode when the position error between the current robot position and root position of reference motion is larger than $1(^{2})$, or the pose error between current robot pose and root pose of reference motion is larger than $\frac{1}{2}\pi$. We set the maximum simulation time for each episode to $20$ which corresponds to maximum episode length = 500.
Reward Coefficients: In our experiment, we use $w_{func}^{ori}=w_{func}^{pos-xy}=0.1875$, $w_{func}^{pos-z}=1.5$, $w_{style}^{adv}=1$, $w_{style}^{joint}=0.5$.
Other Rewards: To smooth the robot trajectory, we also include energy penalty $r_{\text{energy}}$, and action smooth reward $r_{\text{action}}$. $r_{\text{energy}}=-1e-3*\sum_{i}|\tau_{i}\times\dot{q}_{i}|$ where $\tau_{i}$ is the the motor torques applied to the $i$th joint, and the $\dot{q}_{i}$ is the joint velocity for the $i$th joint. $r_{\text{action}}=-1e-2*\sum_{i}|a_{i}^{t}-a_{i}^{t-1}|$ where $a_{i}^{t}$ and $a_{i}^{t-1}$ are the action from policy for $i$th joint at current timestep and the previous timestep.
Simulation Setup: We utilize IsaacGym[32] to simulate 4096 robots in parallel. During motion prior training, for each robot, we uniformly sample a reference motion from the dataset for it to track.
IV Experiments
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Real World Jump Forward Trajectory Comparison: Each row represents a single trajectory (From top to bottom: Reference Motion, VIM, GAIL, Motion Imitation). Trajectories are shown from right to left.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 6: Real World Backflip Trajectory Comparison: Each row represents a single trajectory (From top to bottom: Reference Motion, VIM, GAIL, Motion Imitation). Trajectories are shown from left to right.
We evaluate our system in simulation and real-world, benchmarking it against prior work for low-level locomotion skill learning and various High-level tasks. Our robot demonstrates life-like agility in the real world.
TABLE II: Evaluation of Motion Prior in Simulation: We compare Horizontal and Vertical Root Position (Root Pos (XY), Root Pos (Height)), Root Orientation (Root Ori), Joint Angle, and End Effector Position (EE Pos) tracking errors and RL objectives of all methods. Our methods outperform all baselines in terms of smaller tracking errors, higher episodic returns, and longer episode lengths. GAIL baseline shows a smaller root position tracking error since it can’t follow the reference motion leading to early termination of the episode.
Tracking Error $\downarrow$ RL Objectives $\uparrow$
Method Root Pos (XY) Root Pos (Height) Root Ori Joint Angle EE Pos Episode Return Episode Length
(2) (2) (2) (2) (2)
VIM $\mathbf{1.24\scriptstyle{\pm 0.62}}$ $0.01\scriptstyle{\pm 0.02}$ $0.11\scriptstyle{\pm 0.06}$ $\mathbf{0.08\scriptstyle{\pm 0.06}}$ $\mathbf{0.03\scriptstyle{\pm 0.03}}$ $13.313\scriptstyle{\pm 11.48}$ $166.783\scriptstyle{\pm 120.217}$
VIM (w/o Scheduling) $1.28\scriptstyle{\pm 0.67}$ $0.009\scriptstyle{\pm 0.0123}$ $\mathbf{0.1\scriptstyle{\pm 0.06}}$ $0.1\scriptstyle{\pm 0.08}$ $0.05\scriptstyle{\pm 0.04}$ $\mathbf{13.963\scriptstyle{\pm 11.395}}$ $\mathbf{179.047\scriptstyle{\pm 121.788}}$
Motion Imitation $1.39\scriptstyle{\pm 0.66}$ $\mathbf{0.0077\scriptstyle{\pm 0.0114}}$ $0.11\scriptstyle{\pm 0.05}$ $0.25\scriptstyle{\pm 0.14}$ $0.14\scriptstyle{\pm 0.08}$ $9.536\scriptstyle{\pm 9.049}$ $143.393\scriptstyle{\pm 114.514}$
GAIL $1.04\scriptstyle{\pm 0.86}$ $0.03\scriptstyle{\pm 0.03}$ $0.13\scriptstyle{\pm 0.05}$ $0.17\scriptstyle{\pm 0.1}$ $0.09\scriptstyle{\pm 0.05}$ $3.586\scriptstyle{\pm 6.166}$ $54.723\scriptstyle{\pm 75.984}$
WASABI $0.54\scriptstyle{\pm 0.68}$ $0.03\scriptstyle{\pm 0.03}$ $0.13\scriptstyle{\pm 0.06}$ $4.14\scriptstyle{\pm 1.06}$ $0.21\scriptstyle{\pm 0.07}$ $0.71\scriptstyle{\pm 0.58}$ $22.82\scriptstyle{\pm 16.13}$
TABLE III: Evaluation of Motion Prior in Real (Left): We collect representative metrics for different locomotion skills with corresponding metrics from reference motion. $N/A$ denotes completely failed skills in real.
Metrics Unit VIM Motion Imitation GAIL Reference Motion
Height (Jump While Running) $(m)$ $\mathbf{0.50\scriptstyle{\pm 0.003}}$ $0.42\scriptstyle{\pm 0.01}$ $0.41\scriptstyle{\pm 0.04}$ $0.53\scriptstyle{\pm 0.005}$
Height (Jump Forward) $(m)$ $\mathbf{0.44\scriptstyle{\pm 0.01}}$ $0.42\scriptstyle{\pm 0.01}$ $0.27\scriptstyle{\pm 0.006}$ $0.59\scriptstyle{\pm 0.006}$
Height (Jump Forward (Syn)) $(m)$ $\mathbf{0.52\scriptstyle{\pm 0.01}}$ $N/A$ $N/A$ $0.55\scriptstyle{\pm 0.007}$
Height (Backflip) $(m)$ $\mathbf{0.62\scriptstyle{\pm 0.01}}$ $0.49\scriptstyle{\pm 0.01}$ $N/A$ $0.60\scriptstyle{\pm 0.005}$
Distance (Jump While Running) $(m)$ $\mathbf{0.48\scriptstyle{\pm 0.08}}$ $0.35\scriptstyle{\pm 0.02}$ $0.40\scriptstyle{\pm 0.003}$ $0.56\scriptstyle{\pm 0.008}$
Distance (Jump Forward) $(m)$ $\mathbf{0.76\scriptstyle{\pm 0.05}}$ $0.40\scriptstyle{\pm 0.01}$ $0.10\scriptstyle{\pm 0.002}$ $0.82\scriptstyle{\pm 0.003}$
Distance (Jump Forward (Syn)) $(m)$ $\mathbf{0.49\scriptstyle{\pm 0.04}}$ $N/A$ $N/A$ $0.54\scriptstyle{\pm 0.007}$
Linear Velocity (Pace) $(m/s)$ $\mathbf{0.76\scriptstyle{\pm 0.01}}$ $0.97\scriptstyle{\pm 0.07}$ $0.50\scriptstyle{\pm 0.02}$ $0.72\scriptstyle{\pm 0.05}$
Linear Velocity (Canter) $(m/s)$ $\mathbf{1.49\scriptstyle{\pm 0.15}}$ $N/A$ $N/A$ $3.87\scriptstyle{\pm 0.17}$
Linear Velocity (Walk) $(m/s)$ $0.90\scriptstyle{\pm 0.04}$ $\mathbf{0.96\scriptstyle{\pm 0.06}}$ $0.53\scriptstyle{\pm 0.58}$ $0.97\scriptstyle{\pm 0.42}$
Linear Velocity (Trot) $(m/s)$ $1.33\scriptstyle{\pm 0.17}$ $\mathbf{1.05\scriptstyle{\pm 0.02}}$ $0.93\scriptstyle{\pm 0.01}$ $1.16\scriptstyle{\pm 0.12}$
Angular Velocity (Left Turn) $(rad/s)$ $1.71\scriptstyle{\pm 0.04}$ $0.00\scriptstyle{\pm 0.00}$ $\mathbf{0.91\scriptstyle{\pm 0.04}}$ $1.01\scriptstyle{\pm 0.05}$
Angular Velocity (Right Turn) $(rad/s)$ $0.81\scriptstyle{\pm 0.02}$ $\mathbf{0.62\scriptstyle{\pm 0.02}}$ $0.63\scriptstyle{\pm 0.05}$ $0.41\scriptstyle{\pm 0.09}$
Joint Angle Tracking Error ($rad^{2}/\textit{joint}$) $\mathbf{0.10\scriptstyle{\pm 0.08}}$ $0.27\scriptstyle{\pm 0.16}$ $0.22\scriptstyle{\pm 0.10}$
Figure 7: Latent Skill Space t-SNE. We visualize the latent skill space by visualizing the latent embedding for varying motion segments.
Figure 8: High-level Tasks Evaluation in Simulation: Solid line and shaded area denote the mean and std across random seeds. Our system outperforms all baselines on all tasks.
IV-A Evaluation of Learned Motion Priors
Our system’s proficiency in learning a range of agile locomotion skills from the reference motion dataset (discussed in Sec III-A) is initially assessed.
Baselines: We benchmark our method against two representative baselines: Motion Imitation [40, 5, 22] baseline represents a thread of recent works whose imitation rewards are defined solely with errors between current robot states and the corresponding reference states. Generative Adversarial Imitation Learning (GAIL) baseline represents a thread of recent work [42], whose imitation reward is solely provided by the discriminator trained to distinguish trajectories generated by the policy from the ground truth reference motions. WASABI baseline represents a modified version of WASABI [30] for our setting. Given that our reference motions consist only of state sequences, they offer less supervision compared to expert action sequences, rendering motion prior learning more challenging. Each method trains for $2\times 10^{9}$ iterations across 3 random seeds. Both our method and the Motion Imitation baseline adopt identical reward scales for all motion error-tracking rewards. Likewise, our approach and GAIL maintain the same scale for adversarial stylization rewards.
Figure 9: Real World high-level Following Commands trajectory: Our high-level Following Command policy can track wide-range linear and angular velocity commands even for velocities absent in the reference motion dataset, indicating high-level policy can manipulate the motion prior for High-level tasks. The trajectory is shown from left to right, from top to down.
TABLE IV: High-level Tasks in Real World: We compare Following Command + Jump Forward policies of all methods in real. $N/A$ denotes completely failed skills in real. Our methods outperform all baselines in real for most metrics.
Metrics Unit Ours AMP PPO HRL
Max Linear Velocity $(m/s)$ $\mathbf{1.78\scriptstyle{\pm 0.13}}$ $1.74\scriptstyle{\pm 0.21}$ $1.75\scriptstyle{\pm 0.26}$ $1.70\scriptstyle{\pm 0.08}$
Max Angular Velocity (Left) $(rad/s)$ $1.78\scriptstyle{\pm 0.004}$ $1.07\scriptstyle{\pm 0.09}$ $\mathbf{2.24\scriptstyle{\pm 0.05}}$ $0.00\scriptstyle{\pm 0.00}$
Max Angular Velocity (Right) $(rad/s)$ $\mathbf{2.05\scriptstyle{\pm 0.02}}$ $0.83\scriptstyle{\pm 0.09}$ $1.75\scriptstyle{\pm 0.19}$ $0.95\scriptstyle{\pm 0.37}$
Jump Distance $(m)$ $\mathbf{0.50\scriptstyle{\pm 0.07}}$ $0.00\scriptstyle{\pm 0.00}$ $N/A$ $N/A$
Jump Height $(m)$ $\mathbf{0.50\scriptstyle{\pm 0.02}}$ $0.38\scriptstyle{\pm 0.01}$ $N/A$ $N/A$
Simulation Evaluation: In the simulation, we measure average imitation tracking errors for various agile locomotion skills, episode returns, and trajectory lengths across random seeds. Specifically as listed in Table II, where the tracking error of root pose represents the ability of the robot to reproduce the locomotion skill, and the tracking error of joint angle and end effector position represents the ability of the robot to mimic the style of reference motion. Our method achieves a similar root pose tracking error as the motion imitation baseline with a much smaller joint angle tracking error. This shows our method striking a balance between functionality and style, superior to the motion imitation baseline that focuses solely on functionality. Meanwhile, the GAIL baseline failed to learn the functionality of the reference motions which leads to short episode length and the least episode return. We surmise that the GAIL baseline’s inadequacy arises for two main reasons: First, exclusive reliance on adversarial stylization reward does not offer temporally consistent guidance throughout skill learning due to misaligned rewards across timesteps. Second, the mode collapse issue inherent in adversarial training hinders the robot from mastering highly agile skills, such as backflipping. The shortcomings of the Motion Imitation baseline may stem from the challenges of balancing different terms and selecting suitable hyperparameters when concurrently learning multiple agile locomotion skills. Comparing our VIM with and without stylization reward scheduling, we find the former exhibits enhanced style tracking performance, underscoring the value of stylization reward scheduling in refining robot gait tracking.
Real World Evaluation: We gauge learned agile locomotion skills in real-world scenarios. Due to the lack of precise robot pose, we resort to specific metrics tailored to different locomotion skills, detailed in Table III. For Jump While Running & Jump Forward & Jump Forward (Syn) & Backflip, we measure the jumping height and jumping distance. For Pace & Canter & Walk & Trot and Left Turn & Right Turn, we measure the linear and angular velocity, respectively. Results reveal that our method retains most of the reference motion functionality. The only significant deviation, observed in the Canter motion, arises from inherent differences between animal movement (its source) and our robot’s capabilities since quadrupeds could utilize tendons at their ankle to achieve higher running speed with the specified motion which is not possible for our robot. Even with comparable root pose tracking errors in simulations, our method outshines the Motion Imitation baseline in real-world metrics like jumping height, distance, and velocity tracking error. This suggests that mirroring the style of the reference motion improves sim2real transfer for natural gaits. GAIL baseline struggled to reproduce most real-world locomotion skills. A visual comparison of real-world trajectories is available in Figure  5 and Figure 6, showing our method’s superiority in capturing both motion functionality and style. Due to the poor performance of WASABI baseline in simulation, we didn’t evaluate it in the real world.
Latent Skill Space Visualization: To better understand our low-level motion prior, we visualize the learned latent skill space in Figure 8 by visualizing the latent embedding corresponding to motion segments in our reference motion dataset via t-SNE [53]. We find that different skills are separated into different regions with clear boundaries. Our reference motion encoder also clusters the skills with similar semantic meaning together: the embeddings from Left Turn sequence are close to the embeddings from Right Turn sequence, which enables the smooth transition between different skills, the embeddings from Jump While Running & Jump Forward & Jump Forward (Syn) sequence are clustered together. These observations suggest that our system learned a smooth and semantically meaningful latent skill space for solving High-level tasks.
TABLE V: Ablation Study of Motion Prior in Simulation: We compare our method and two variant of our method without Functionality (Func) reward or Style reward on all different objectives
Tracking Error $\downarrow$ RL Objectives $\uparrow$
Method Root Pos (XY) Root Pos (Height) Root Ori Joint Angle EE Pos Episode Return Episode Length
(2) (2) (2) (2) (2)
VIM $\mathbf{1.24\scriptstyle{\pm 0.62}}$ $0.01\scriptstyle{\pm 0.02}$ $0.11\scriptstyle{\pm 0.06}$ $\mathbf{0.08\scriptstyle{\pm 0.06}}$ $\mathbf{0.03\scriptstyle{\pm 0.03}}$ $\mathbf{13.313\scriptstyle{\pm 11.48}}$ $\mathbf{166.783\scriptstyle{\pm 120.217}}$
VIM (w/o Func Reward) $1.24\scriptstyle{\pm 0.67}$ $0.01\scriptstyle{\pm 0.02}$ $0.11\scriptstyle{\pm 0.06}$ $0.61\scriptstyle{\pm 0.59}$ $0.02\scriptstyle{\pm 0.02}$ $10.66\scriptstyle{\pm 12.92}$ $115.19\scriptstyle{\pm 117.74}$
VIM (w/o Style Reward) $1.49\scriptstyle{\pm 0.69}$ $\mathbf{0.00\scriptstyle{\pm 0.01}}$ $0.12\scriptstyle{\pm 0.06}$ $5.14\scriptstyle{\pm 1.71}$ $0.25\scriptstyle{\pm 0.06}$ $6.28\scriptstyle{\pm 6.73}$ $109.76\scriptstyle{\pm 110.67}$
Figure 10: Performance for different reference motions: We provide the average episode length of the learned motion prior when it imitates different reference motions.
IV-B Evaluation on High-level Tasks
To further evaluate how our method leverages learned agile locomotion skills for High-level tasks, we designed a set of tasks and tested our method against standard baselines in both simulation and real-world scenarios.
High-level Tasks: Our task suite comprises: Following Command: This involves directing the robot to move with specific linear and angular velocities. The linear velocity command and angular velocity command are sampled uniformly between $0\sim 2$ m/s and $-2\sim 2$ rads/s respectively. In our motion prior, the robot is trained to move and turn at the reference motion’s speed; hence, to follow a command precisely, the high-level policy should smoothly interpolate between different speeds. Jump Forward: This task requires the robot to jump while running forward. We have adapted a subset of jumping rewards from CAJun [58] to evaluate policy interpolation between jumping and running motions within a fixed timeframe. Following Command + Jump Forward: Here, the robot must either jump forward or adjust to changing commanded speeds. To optimize episode return, the robot should not only use the agile locomotion skills from the reference motion dataset but also develop unobserved skills like executing sharp turns.
High-level Observation For Following Command task, our high-level observation includes the target linear velocity and target angular velocity. For Jump Forward task, our high-level observation includes the target jumping forward velocity and the normalized phase information in the jumping forward cycle. For Following Command + Jump Forward task, our high-level observation includes the high-level observation for both Following Command and Jump Forward tasks as well as an additional binary command indicating whether following command or jump forward at current time-step
Baselines: Considering the baseline’s subpar performance in low-level motion prior training, we compare our system with three representative baselines without pre-trained low-level controllers: PPO [45]: Demonstrates controllers trained exclusively on High-level task rewards. AMP [41] utilize existing reference motion to provide styling reward in an adversarial imitation learning manner and learn the policy for the High-level task while mimicking the behavior of reference motions. Hierarchical Reinforcement Learning (HRL) adapts from Jain et al. [25] which learns a high-level policy sending latent commands to a low-level motor controller. HRL resembles a broad category of prior works that decompose temporally extended reasoning into sub-problems [50, 11, 38, 35, 18, 28, 21]. For a fair comparison, we made modifications like removing the trajectory generator in [25], using PPO for AMP and HRL, and supplying full reference motion data to AMP and HRL integrated with AMP.
Evaluation in Simulation & Real World: We train all methods on each High-level task for $4\times 10^{8}$ environment samples with 3 random seeds. The simulation results are detailed in Figure 8, and real-world results are provided in Table IV. Real-world Following Commands trajectory is also provided in Figure 9. For the Following Command task, while all methods mastered basic locomotion, ours excelled in efficiency and smoothly transitioned between diverse linear and angular velocities. The other tasks, Jump Forward and Following Command + Jump Forward, demanded advanced jumping abilities, which baselines couldn’t emerge. These baseline methods either continuously moved forward, remained grounded when prompted to jump, or toppled to evade energy consumption penalties. In contrast, our system seamlessly bridged jumping and running actions, securing the highest episode return. Despite providing with a comprehensive reference motion dataset, baselines couldn’t harness the skills. This shortcoming possibly stems from the challenge of deriving agile locomotion skills from the dataset using only adversarial stylization rewards, mirroring the GAIL baseline’s poor performance in low-level motion prior training.
IV-C Ablation Study of Learned Motion Prior
To further understand how different reward and reference motions influence our framework, we provide further ablation study over reward terms as well as a more fine-grained performance of our framework when it imitates different reference motions.
When we remove the Functionality (Func) reward or Style reward, we found the performance of our method drops to a different extent, as shown in Table V . We found that without Funtionality reward, the learned controller could not robustly track the reference motion resulting in smaller Episode Return and, shorter Episode Length. On the other hand, removing Style reward results in a significantly higher Joint Angle and End-Effector tracking error.
In our framework, when the root pose of the simulated robot diverges too much from the reference root pose, we terminate the episode, as described in Sec III-E . In this case, the episode length is a good indicator of whether the learned policy could follow the reference motions. As shown in Fig. 10 , the performance of our framework varies when imitating different reference motions. When imitating relatively steady motions like Walk (Mocap), Pace (Mocap), Left Turn (Mocap), the learned controller could track the motion for a longer period. When imitating relatively agile motions, especially with high moving speed, such as Canter (Mocap), Jump While Running (Mocap), the performance of our system drops. This phenomenon is rooted in the methodology disparities between our robot and real animals.
V Conclusion
In this paper, we propose Versatile Instructable Motion prior (VIM) which learns agile locomotion skills from diverse reference motions with a single motion prior. Our results in simulation and in the real world show that our VIM captures both the functionality and the style of locomotion skills from reference motions. Our VIM also provides a temporally consistent and compact latent skill space representing different locomotion skills for different High-level tasks. With agile locomotion skills in our VIM, complex High-level tasks can be solved efficiently with minimum human effort.
Supplementary Material
I Detailed Observation Space
we provide more detailed observation space for our motion prior. Our Unitree A1 robot has 12 joints, corresponding to 12 Degrees of Freedom (DoF), and we use positional control for the 12 DOF ($KP=40$ and $KD=1.0$). Specifically, the proprioceptive state of the robot contains:
  • Joint Angle - $\mathbb{R}^{12\times 3}$ contains joint rotations for all joints ($12$D) for the past three control step.
  • Joint Velocity - $\mathbb{R}^{12\times 3}$ contains joint velocities for all joints ($12$D) for the past three control step.
  • Previous Action - $\mathbb{R}^{12\times 3}$ contains positional command for all joints ($12$D) for the past three control step.
  • Projected Gravity - $\mathbb{R}^{3\times 3}$ contains the projected gravity in the robot frame, representing the orientation of the robot for the past three control steps.
  • Foot position - $\mathbb{R}^{3\times 4\times 3}$ contains the robot foot positions in the robot frame, $3$ dim per foot per timestep for the past three control steps
Note that, our discriminators used for adversarial reward only use the joint angle transition for training and reward calculation.
We also provide additional high-level observation for the example high-level tasks we used. For Following Command task, we provide target linear velocity and target angular velocity as high-level observation. For Jumping Forward task, since the robot is tasked to jump forward in a fixed frequency, we provide normalized temporal phase as high-level observation.
II Domain Randomization
Here we provide our hyperparameters related to domain randomization for better sim2real transfer and shared by all methods.
Parameter Range
Added Mass for the base [-1, 3]
Friction [0.1, 1.3]
Restitution Range [0, 1.0]
COM shift [-0.05, 0.05]
Motor Strength ratio [0.7, 1.1]
KP randomize ratio [0.8, 1.5]
Kd randomize ratio [0.5, 1.5]
Proprioception noise [0, 0.01]
Action noise [0, 0.05]
III RL Training Details
Here we provide hyperparameters related to RL training and shared by all methods.
Hyperparameter Value
Max Episode Length 500
Non-linearity ELU
Policy initialization Standard Gaussian
# of samples per iteration 4096 * 48
Discount factor .99
Parallel Environment 4096
Optimization epochs 5
# of batches 16
Clip parameter 0.1
Policy network learning rate 3e-4
Value network learning rate 3e-4
Discriminator learning rate 1e-5
Entropy 0.001
Optimizer Adam
$\beta$ for latent regularization 1e-5